Meeting 2024 08 07
From Hall A Wiki
GMn GEM commissioning Meetings: WEDNESDAY 10:30am - 11:30am
Link to all previous meetings: GMn GEM Commissioning Meeting
Please use sbs_gems mailing list for plots/communications and HALOG entries relevant to GEMs!
Contents
Progress from this past week
- DSG beginning the movement of the BB gas flowmeters to the SBS line.
- Front tracker is now fully assembled[1].
- First 6 layers (L8-L13) haven been installed into back tracker
- M0 and M2 have been replaced in L14 (SBS L6)[2][3].
- The BB crates have been moved over to SBS rack. We will start planning on how to connect the MPDs.
GEp preparations
- Our GEM task list is here: workplan
- Remove layers, isolate shorted sectors, repair APVs as needed
- Joe B will add LV inputs for new GEMs in the stack, we will need to make those cables (4 more layers)
- Install BB gas distribution from SBS line -- DSG will do this
- Further calibrations of SBS gas mix
- Repair/troubleshoot problematic CAEN crates -- DSG working on this
- MPD testing and repair
- Prepare and install pixel layers (long term)
Suggested order of work
- Move layers into front and back tracker (done)
- Upgrade the BB XY layer to SHV (done)
- Fabricate remaining LV cables
- Repair/replace L6 modules, install as last layer in Back Tracker, until EEL layer is ready (work in progress)
- For cosmics, bring bars (HV+signal cables) to table and make a readout trigger. Work with techs to bring out 1 layer at a time?
- For cosmics, move stack to be near RHRS, we can access all gas lines, setup readout, LV etc. (completed)
- Move BB crates to SBS rack (completed)
- Setup scintillators around the stack and run for a longer period. (work in progress)
- Other work, setup gas flow in EEL to chamber. Finish HV upgrade & test. (work in progress)
DAQ updates (Ben, Bryan, Alex)
- Implement and test histogram method